OpenStudy (anonymous):

In ΔABC shown below, use an appropriate trigonometric ratio to find m∠B. Round to the nearer degree. A. 60 B. 53 C. 37 D. 30

OpenStudy (anonymous):



OpenStudy (anonymous):

cos b = adj/hyp cos b = 4 / 5

OpenStudy (anonymous):

SOH CAH TOA nice trick also if you understand it

OpenStudy (anonymous):

cos b = 80 degrees but that is not one of your choices... using pyth theorem the other side is 3 (special right triangle 3-4-5) sin b = 3/5 = 60 degrees. not sure why cos did not work...let me think on that (they should have both came out to the same degrees